OpenStudy (anonymous):

Which sequence is equivalent to t(n) = 3 + 6n, n > 1 where ? Answer a. t(1) = 9; t(n + 1) = 6t(n) b. t(1) = 9; t(n + 1) = t(n) + 6 c. t(1) = 3; t(n + 1) = t(n) + 6 d. t(1) = 3; t(n + 1) = 6t(n)

OpenStudy (anonymous):

t(1) = 3+6(1) = 9 so it is choice a or b. Then use t(1) to see what t(1+1) is using the original problem t(2)=3+6(2) = 15 see which choice a or b gives you 15 when you use 1 for n

OpenStudy (amoodarya):

b. t(1) = 9; t(n + 1) = t(n) + 6 put 1 you have 9 put 2 you have 15 put 3 you have 21 ... 9,15,21,27, ... then t(n + 1) - t(n) = 6

OpenStudy (anonymous):

b is correct, nice job